OpenStudy (anonymous):

Write the equation in standard form for the parabola with vertex (0, 0), and focus (0, 1).

OpenStudy (anonymous):

I got y=1/2 which isn't even on the answer option list

OpenStudy (anonymous):

It's asking for the equation. o.o

OpenStudy (anonymous):

The answer options are: 1.)x=1/2y^2 2).4y^2 3)y=4x^2 4)y=1/2x^2

OpenStudy (anonymous):

I know it is asking for the equation

OpenStudy (kohai):

\[(x-h)^2 = 4p(y-k)\] (h, k) as the vertex p = distance from the vertex to the focus p > 0, opens upward p < 0, opens downward

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

They gave you the equation. Now all you have to do is plug the numbers you gave us into it.

OpenStudy (anonymous):

Hhaha ok that's not the equation I was taught

OpenStudy (anonymous):

And I don't know how to plug it in to that I've never seen it before

OpenStudy (kohai):

What equation were you taught?

OpenStudy (anonymous):

the distance formula

OpenStudy (kohai):

I'm not sure how you would use the distance formula for this, because you would just get d = some value

OpenStudy (kohai):

And all of the sites I'm looking at for this kind of question uses the formula I gave

OpenStudy (anonymous):

I don't know I'm really confused

OpenStudy (anonymous):

You don't use the distance formula to get an equation for a parabola. O.o

OpenStudy (anonymous):

Sorry but my textbook does

OpenStudy (kohai):

Can you do me a favor and take a picture of the example your textbook does for this problem? So I can see what you're looking at?

OpenStudy (anonymous):

√(x-x1)^2+(y-y1)^2=√(x-x2)^2+(y-y^2)^2

OpenStudy (kohai):

Oh, I see. \[\sqrt{(x-x1)^2+(y-y1)^2} = \sqrt{(x-x2)^2+(y-y2)^2}\]